I need to write and equation in slope -intercept form of the line that passes through the given point and is parrallel to the graph of each equation. 2x + y =5;(3,1). Please help.
Thank you
cathy

Answer by algebrapro18(249)

well first we need to solve the given equation for y

2x+y=5
y=-2x+5

we now know that the slope is -2 and since parallel lines have the same slope we know the line we are getting the equation to has a slop of -2. From here there are 2 ways to do this problem we can either use point slope formula or sub in the point and slope into the eqation y=mx+b and solve for b.

  • Using point slope formula

    y-y1=m(x-x1)

    where y1=1, x1=3, and m = -2

    y-1=-2(x-3)
    y=-2x+6+1
    y=-2x+7
  • using algebra

    y=mx+b

    where y=1 m=-2 and x=3

    1=-2(3)+b
    1=-6+b
    7=b

    y=-2x+7